OpenStudy (anonymous):

the formula for the volume of a sphere is V=4/3pir^3. find the radius, to the nearest hundredth, of a sphere with a volume of 15in^3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\huge V=\frac{4}{3}\pi r^3$$

OpenStudy (anonymous):

no i dont im very confused

OpenStudy (skullpatrol):

The volume is given as $$15 in^3$$ and the question asks you to find the radius...

OpenStudy (anonymous):

how would you find the radius though

OpenStudy (skullpatrol):

Solve the formula for it.

OpenStudy (skullpatrol):

$$\huge 15=\frac{4}{3}\pi r^3$$

OpenStudy (anonymous):

oh okay so do you like multiply it to both sides or osmething

OpenStudy (skullpatrol):

Yes, use the multiplication property of equality...

OpenStudy (skullpatrol):

$$\huge \frac{15}{\frac{4}{3}\pi}= r^3$$

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so then you just divid 15 by 4/3 right ?

OpenStudy (skullpatrol):

$$\pi \approx 3.14$$

OpenStudy (anonymous):

so you times 3.14 by 4/3 correct ?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

okay so that would be 4.186 so then \[\frac{ 15 }{ 4.186}=r ^{3}\]

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

so then you divide 15 by 4.186 right

OpenStudy (skullpatrol):

right

OpenStudy (anonymous):

okay and then youll get 62.79=r^2

OpenStudy (anonymous):

so is that the answer ?

OpenStudy (skullpatrol):

$$\huge r^3=62.79$$

OpenStudy (skullpatrol):

$$\huge\text{" find the radius,r"}$$ $$\huge\text{" to the nearest hundredth,"}$$

OpenStudy (skullpatrol):

$$\huge (r^3)^{\frac{1}{3}}=(62.79)^{\frac{1}{3}}$$

OpenStudy (anonymous):

so r= 18.837

OpenStudy (skullpatrol):

to the the nearest hundredths or 0.01 and don't forget the units...

OpenStudy (anonymous):

so it would be 19 ?

OpenStudy (skullpatrol):

Nope, are you using a calculator?

OpenStudy (anonymous):

yeah i used my phone

OpenStudy (skullpatrol):

Find $$\huge(62.79)^{0.333}$$

OpenStudy (anonymous):

so it would be 20.900907

OpenStudy (skullpatrol):

Nope, that is too large.

OpenStudy (anonymous):

well thats what my calculator says....

OpenStudy (skullpatrol):

3.97 inches

OpenStudy (anonymous):

is the answer ?

OpenStudy (skullpatrol):

Yes, now check the answer in the original equation.

OpenStudy (skullpatrol):

$$\huge V=\frac{4}{3}\pi r^3$$ $$\huge 15=?=\frac{4}{3}\pi (3.97)^3$$

OpenStudy (skullpatrol):

r= 1.53 in